User Tools

Site Tools


handover:aua026

aua026

Hobart 26m:

Disk VSN: HOB+0098

Data Volume at Beginning: 37.50 GB

  • 18:00 Experiment started OK (Ross)
  • Maser has been synced but reports a syncerr_gt_3 which can not be corrected in fmset. Jamie says maserdelay is from different maser (Ross)
  • 23:42 Mark 5B syncerr_gt_3 error, ran fmset on hobart but not sure if it will hold. dbbc=pps_delay was jumping around a bit previously, will monitor it. (Patrick)
  • 04:17 Weather command missing from midob procedure, added in by Jamie. (Patrick)
  • 17:06 He supply pres has dropped to approx 273psi but is climbing again. A similar large drop in pressure then increase occurred between approx 1200-1500 but not so low. Now above 275psi again… (Liz)

Katherine 12m:

Disk VSN: HOB+0068

Data Volume at Beginning: 0.00 GB

  • 18:00 Experiment started OK (Ross)
  • GPIB has been removed so no clkoff or maserdelay measurements reported (Ross)
  • 23:36 Ke reporting overflowing Tsys measurements in all channels, and autocorrelation spectra looks super weird. Power cycled DBBC, reconfigured, and synched times to fix it. Resumed schedule at 00:10 with scan No0082 (Patrick)
  • 00:33 Overflowing Tsys measurements again. Restarted and reconfigured DBBC as before. Resumed at 00:47 with scan No0093. (Patrick)
  • 05:10 Tsys overflowing again, restarted and reconfigured DBBC as before. mk5 wouldn't resync, got this error when running fmset: mk5cn: time-out, connection closed. Tried restarting Mark5 (Dimino process), and reconfiguring DBBC again, but no luck. Jamie looking in to it. (Patrick)
  • 06:00 DBBC needs fixing onsite, a card may need replacing. Katherine is out of the experiment, and hopefully will be fixed tomorrow. (Patrick)
  • 07:20 Mick fixed issues with DBBC, good data again at scan no0164. (Bryn)
  • 10:00 network issue, lost connection to mk5ke. getting error: no route to host. power cycling mk5 and reformatting time with fmset. Missed scan no0189 and no0190. (Bryn)

Yarragadee 12m:

Disk VSN: GSFC+015

Data Volume at Beginning: 0.00 GB

  • 18:00 Experiment started OK (Ross)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/aua026.txt · Last modified: 2017/09/14 17:08 by Jonathan Rogers